OpenStudy (anonymous):

A solution to the equation x2 – 9x – 22 = 0?

OpenStudy (lgbasallote):

do you know how to factor out \(x^2 - 9x -22\)?

OpenStudy (anonymous):

Yes. Is the answer -11?

OpenStudy (lgbasallote):

no..what was your factored form?

OpenStudy (anonymous):

9x+2(x)+1=0 (To solve for x) x= - 1/11. X= negative one over eleven.

OpenStudy (anonymous):

\[x^2 – 9x – 22 = 0\] \[d=b^2-4ac\] \[d=(9)^2-4(1)(-22)\] \[d=81+88=169\] Therefore, \[x=(-b \pm \sqrt{b^2-4ac})/2a\] Putting the valuen we get, x=11 or x=-2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Welcome ;)